OpenStudy (harley1312):

Help please, I will give medal and fan... M=1,(-2,3) write the equation of the line in point slope form and in Slope intercept form

OpenStudy (princeharryyy):

y = 1x + c 3 = 1*(-2) +c c =5 equation >> y =x +5

OpenStudy (triciaal):